OpenStudy (anonymous):

What is the y coordinate of the y-intercept of the line that passes through the points (-4, -1) and (4, 5)?

OpenStudy (anonymous):

put it in y=mx+b form where b is your answer and m is your slope. slope is (5+1)/(0) hmm this immediately tells me there is a problem or no solution cuz of division by zero. or i did something wrong lol.

OpenStudy (anonymous):

lets draw to see maybe the coordinants are just in wrong order  yeah just in wrong order

OpenStudy (anonymous):

back y =mx+b m=(-1-5)/(-4-4)=.75 so now we chose a point that is given by your coordinants to solve for b 5=.75*4+b sfb b=2 that is your answer 2